Hello! I'm having trouble with figuring out slope intercept when y is a fraction. For example, 2/5y=6x-2
Answer by mananth(16946)

(2/5)y=6x-2
Since the 5 in the denominator is a bother remove it
Multiply the equation by 5

(2/5)*5 y= 6*5x-2*5
2y=30x-10
/2
y=15x-5
m=15
and intercept = -5